### v1

# Language Integration in Fine-Tuning Multimodal Large Language Models for Image-Based Regression

Abstract
Multimodal Large Language Models (MLLMs) show promise for image-based regression tasks, but current approaches face key limitations. Recent methods fine-tune MLLMs using preset output vocabularies and generic task-level prompts (e.g., ”How would you rate this image?”), assuming this mimics human rating behavior. Our analysis reveals these approaches provide no benefit over image-only training. Models using preset vocabularies and generic prompts perform equivalently to image-only models, failing to leverage semantic understanding from textual input. We propose Regression via Transformer-Based Classification (RvTC), which replaces vocabulary-constrained classification with a flexible bin-based approach. Unlike approaches that address discretization errors through complex distributional modeling, RvTC eliminates manual vocabulary crafting through straightforward bin increase, achieving state-of-the-art performance on four image assessment datasets using only images. More importantly, we demonstrate that data-specific prompts dramatically improve performance. Unlike generic task descriptions, prompts containing semantic information about specific images enable MLLMs to leverage cross-modal understanding. On the AVA dataset, adding challenge titles to prompts improves correlations from 0.83 to 0.90, a new state-of-the-art. We demonstrate through empirical evidence from the AVA and AGIQA-3k datasets that MLLMs benefit from semantic prompt information surpassing mere statistical biases. This underscores the importance of incorporating meaningful textual context in multimodal regression tasks.
1 Introduction
Vision-language models trained on massive unlabeled image-language datasets have demonstrated remarkable capacity to extract universal image features, with CLIP [15] achieving impressive zero-shot classification performance on benchmark datasets such as ImageNet. Building on these foundations, Multimodal Large Language Models (MLLMs) have evolved to seamlessly fuse image and text embeddings with generative language capabilities [5, 11, 22]. This has sparked growing interest in transferring MLLM capabilities to image-based regression tasks, including Image Quality Assessment (IQA), Image Aesthetics Assessment (IAA) [4, 20], and AI-Generated Image Quality Assessment (AIGIQA) [21, 14].
Figure 1: Existing MLLM methods using preset vocabulary and generic prompts (left) achieve 0.82 correlation on AVA. Our image-only RvTC model (center) exceeds this with state-of-the-art correlation of 0.83. Integrating data-specific prompts (e.g., ”Outdoor Macro Shot”) during fine-tuning (right) unlocks MLLM cross-modal reasoning, yielding a new state-of-the-art 0.90.
Recent approaches [8, 7, 20, 19] fine-tune MLLMs for image regression using two key assumptions borrowed from human rating behavior: (1) utilizing preset output vocabularies (e.g., ”excellent”, ”good”, ”fair”, ”poor”, ”bad”), and (2) incorporating generic task-level prompts such as ”How would you rate the quality of this image?”. These methods assume that mimicking human-like vocabulary and task descriptions will leverage the multimodal capabilities of MLLMs.
However, we demonstrate that current multimodal approaches provide no benefit over image-only training. Models using preset vocabularies and generic prompts perform equivalently to image-only models. This challenges the core assumption that current MLLM fine-tuning strategies effectively utilize cross-modal capabilities for regression tasks. To address these limitations, we make three key contributions:
Table 1: Comparison of models on the AVA dataset. RvTC achieves state-of-the-art correlation of 0.83 without using any textual input. When data-specific prompts (challenge titles) are added (RvTC+), performance increases to 0.90. “LP” refers to linear probing of the regression head; “+” indicates inclusion of challenge titles in prompts.
Method SRCC PLCC
RvTC-LP 0.709 0.711
RvTC-LP+ 0.742 0.741
NIMA [16] 0.612 0.636
MUSIQ [3] 0.726 0.738
VILA [4] 0.774 0.774
LIQE [24] 0.776 0.763
One-Align [20] 0.823 0.819
RvTC (ours) 0.833 0.831
RvTC+ (ours) 0.899 0.901
1. Rethinking regression with RvTC: We propose Regression via Transformer-Based Classification (RvTC), which replaces the rigid vocabulary constraints of previous methods with a flexible bin-based regression scheme. This simple yet effective approach outperforms prior multimodal methods using only images, matching or setting new state-of-the-art on four benchmarks. Unlike recent work that addresses discretization errors through complex distributional modeling [23], RvTC achieves superior accuracy by simply increasing the number of classification bins.
2. Enhancing MLLM performance through data-specific prompts: We demonstrate that generic task prompts (e.g., ”How would you rate this image?”) fail to leverage cross-modal capabilities. In contrast, fine-tuning with data-specific prompts yields substantial improvements in correlation results. For example, the prominent IAA dataset AVA includes challenge titles that characterize images with descriptive phrases such as ”Rule of Thirds” or ”Outdoor Macro Shot.” Incorporating these semantic descriptors during fine-tuning improves correlation results from our already state-of-the-art 0.83 to 0.90.
3. Disentangling semantics from bias: Through controlled experiments on AVA and AGIQA-3k datasets, we show that a significant portion of the observed gains stem from cross-modal understanding rather than dataset-specific statistical artifacts.
Our findings demonstrate that effective fine-tuning of MLLMs for regression tasks requires moving beyond human-like vocabulary and generic prompts. Instead, training should incorporate semantically meaningful, data-specific context that aligns with the model’s cross-modal capabilities. When fine-tuned with such input, MLLMs exhibit significant gains in regression accuracy, revealing their potential for grounded visual understanding.
2 Related work
We highlight the following notable prior works that inform our method.
REgression using CLAssification (RECLA). RECLA involves transforming a regression problem into a classification problem, leveraging the strengths of classification algorithms for predictive tasks [17]. In RECLA, the continuous target variable is discretized into a set of classes, also referred to as bins, effectively categorizing the range of possible values. A classification model is then trained to predict the appropriate bin for each input, and the prediction is often mapped back to a continuous value by a weighted average of the bins’ midpoints. In this work, we show that RECLA seamlessly integrates with pre-trained Multimodal Large Language Models (MLLMs). This integration is robust and a randomly initialized RECLA head can be fine-tuned together with the MLLM without any special modification to the downstream training procedure.
Multimodal Large Language Models. Vision-language models extend the capabilities of traditional language models to incorporate and process multiple modalities, such as image, audio, and video. A key component of Vision-language models is their ability to align representations across modalities, allowing them to understand the relationships between text and other forms of data.
MLLMs, such as LLaVA [11] and mPLUG-Owl2 [22], further enhance this capability by seamlessly fusing image and text embeddings, enabling more complex multimodal reasoning and generation tasks. The architecture of these models involves a combination of modality-specific encoders and a shared transformer-based network that processes the combined representations. The transformer architecture allows the model to capture long-range dependencies and complex interactions between modalities.
Image-based regression tasks using MLLMs. The application of MLLMs to image-based regression tasks has garnered increasing attention, driven by the potential to leverage the rich representations learned from large-scale multimodal datasets. Image Quality Assessment (IQA), Image Aesthetic Assessment (IAA) and AI-Generated Image Quality Assessment (AIGIQA) are prominent areas where MLLMs have shown promising results. In IQA, the goal is to predict the perceived quality of an image, often by learning to map image features to subjective quality scores. In IAA, the objective is to assess the aesthetic appeal of an image, typically by predicting scores that reflect human aesthetic preferences. In AIGIQA, generated images are evaluated for perceptual quality and alignment with the generation prompt.
One approach to improve IAA representations is through specialized pre-training, as demonstrated by VILA [4]. VILA employs a vision-language pre-training strategy to learn representations that are specifically tailored for aesthetic assessment. This involves training the model on a large dataset of images and associated aesthetic scores, using a combination of contrastive learning and regression objectives. By pre-training on aesthetic-specific data, VILA is able to learn representations that are more effective for predicting aesthetic scores compared to general-purpose vision-language models. The recent Q-Align [20] teaches an MLLM for visual rating aligned with human opinions. Q-Align achieves state-of-the-art performance on image quality assessment (IQA), image aesthetic assessment (IAA), as well as video quality assessment (VQA) tasks. Q-Align unifies the three tasks into one model they call OneAlign.
Architecturally, our work generalizes and extends the Q-Align framework by substituting the vocabulary-based classification head of mPLUG-Owl2 with regression using a classification head, resulting in improved performance and broader applicability. In addition, we present and analyze performance gains that are obtained by the incorporation of textual prompts with semantic relevance to the input images during fine-tuning MLLMs for image-based regression tasks.
A concurrent approach, DeQA-Score [23], addresses the limitations of discretizing continuous quality scores into one-hot labels, which can lead to information loss. DeQA-Score proposes to discretize the entire score distribution into soft labels rather than hard classification targets, aiming to preserve more information about the continuous nature of quality scores. While this approach tackles the discretization problem through distributional modeling, our work demonstrates that simply increasing the number of bins in a straightforward regression using classification framework achieves state-of-the-art performance gains without requiring complex distributional representations (see Fig. 3). This suggests that the discretization accuracy problem can be effectively addressed through increased granularity rather than sophisticated label representations.
3 Method
In this section, we present Regression via Transformer-Based Classification (RvTC), a framework that transforms multimodal regression into a classification problem with flexible bin counts. Unlike existing approaches that constrain outputs to preset vocabularies, RvTC eliminates manual vocabulary crafting while achieving superior performance through straightforward bin increase.
3.1 Architecture overview
RvTC builds upon the Multimodal Large Language Model mPLUG-Owl2 [22], which demonstrates strong visual perception and language understanding capabilities. The base architecture comprises: (1) Vision encoder: ViT-L/14 [15] processes input images. (2) Visual abstractor: Reduces visual features to 64 semantic token embeddings per image. (3) Language decoder: LLaMA-2-7B [18] serves as a universal interface for mixed vision-language input. We replace mPLUG-Owl2’s vocabulary-constrained classification head with a $K$-bin linear classification head that supports arbitrary bin counts for regression tasks. The bin classification head is applied to the penultimate hidden-state embedding of the final token that acts as an aggregator for all of the tokens.
3.2 Regression using classification framework
Problem formulation. Building on the RECLA (REgression using CLAssification) framework established by [17], RvTC reformulates regression problems $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ as classification problems $g:\mathbb{R}^{d}\rightarrow{1,2,\ldots,K}$, where $K$ represents the number of bins. This transformation involves: (1) Discretization: Target values are discretized into $K$ distinct bins using uniform binning over a preset min-max range. (2) Assignment: Each target value is assigned to the bin whose center is closest to the original value. (3) Classification: Inputs are classified using a linear head with $K$-bins as classes.
3.3 Training and inference
For training, we use standard cross-entropy loss to optimize bin classification. During inference, posterior probabilities ${p_{1},p_{2},\ldots,p_{K}}$ are computed via softmax, then converted to continuous values through a weighted sum: $\sum_{i=1}^{K}p_{i}b_{i}$, where $b_{i}$ is the center of bin $i$.
3.4 Advantages of the bin-based approach
This formulation offers simplicity and flexibility over vocabulary-constrained methods; bin count can be adjusted without redefining vocabularies, and performance improves monotonically with increased number of bins (see Fig. 3) in contrast to complex distributional modeling [23].
Note that when quantizing the target values of the data, there is freedom in setting the quantization scheme, for example, using a non-uniform quantization scheme. However, increasing the number of bins is a straightforward way to reduce quantization noise and simplify training, removing the need to tune hyperparameters of the range at the cost of added complexity to the classification task.
3.5 Training configurations
Image-only training. When fine-tuning without textual prompts, we refer to the model as image-only RvTC. This configuration serves as our baseline and demonstrates that effective regression can be achieved using only visual features.
Multimodal training with data-specific prompts. For multimodal training, we incorporate data-specific prompts that contain semantic information relevant to individual images, rather than generic task descriptions. This approach enables the model to leverage cross-modal understanding for improved regression performance.
4 Experiments
4.1 Experimental setup
Datasets. We conduct experiments on five datasets spanning different image assessment tasks. For Image Aesthetic Assessment (IAA), we use AVA [13], a prominent benchmark extracted from the DPChallenge website containing over 250,000 photographic images with mean opinion scores (MOS) from 1-10. Each image belongs to one of 1,400 challenges with descriptive titles such as ”Rule of Thirds” and ”Self Portrait Without People”; these challenge titles serve as our data-specific prompts.
For Image Quality Assessment (IQA), we evaluate on three datasets: KonIQ-10k [2] (10k images from the extensive YFCC100M multimedia database), SPAQ [1] (11k smartphone photos), and KADID-10k [10] (10k images with 25 distortion types at 5 intensity levels). For AI-Generated Image Quality Assessment (AIGIQA), we use AGIQA-3k [9], containing 3k AI-generated images with corresponding generation prompts and MOS ratings for semantic alignment and perceptual quality.
All evaluations use official test sets with Spearman’s Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC) as metrics.
Model architecture and training. Our RvTC framework builds on mPLUG-Owl2 [22], using ViT-L/14 [15] as the vision encoder and LLaMA-2-7B [18] as the language decoder. We replace mPLUG-Owl2’s vocabulary classification head with a randomly initialized linear classification head for $K$-bin classification.
We fine-tune the entire model using Adam optimizer [6] with cosine learning rate scheduling [12]. The learning rate is initialized to 1e-5 with 0.03% warm-up steps. For experiments without textual prompts, we use batch size 128 for 2 epochs; with textual prompts, we train for 3 epochs. All experiments use 4 NVIDIA RTX H100 GPUs.
Based on the results presented in Fig. 3, we set the number of bins to 51 and use uniform binning over the preset min-max range of each training dataset, with target values replaced by the bins whose center is closest to the target value.
Dataset-specific adaptations. For AGIQA-3k, due to its smaller size, we first pre-fine-tune on AVA for 2 epochs before task-specific fine-tuning. This transfer learning approach leverages the larger aesthetic assessment dataset to improve performance on the AI-generated image task.
4.2 Baseline performance: image-only RvTC
In this section, we establish RvTC’s image-only baseline performance to isolate the contribution of textual prompts evaluated in Sec. 4.3. Our RvTC approach achieves state-of-the-art results using only visual input across multiple datasets.
Linear probing analysis. Tab. 1 shows results for RvTC-LP (linear probing), where only the regression head is fine-tuned while the backbone remains frozen. This approach achieves strong image-only performance on AVA (SRCC: 0.709, PLCC: 0.711), demonstrating the generalization capabilities of mPLUG-Owl2’s pre-trained visual representations without requiring full model fine-tuning (we refer to RvTC-LP+ and RvTC+ in Sec. 4.3).
Table 2: Performance comparison in SRCC/PLCC of image-only RvTC with different models on IQA tasks
KonIQ-10k SPAQ KADID-10k
NIMA [16] 0.86/0.90 0.91/0.91 NA/NA
MUSIQ [3] 0.93/0.91 0.92/0.92 NA/NA
LIQE [24] 0.92/0.91 0.92/0.92 0.93/0.93
One-Align [20] 0.94/0.95 0.93/0.93 0.94/0.94
[…]
Comparison with existing methods. Comparing with established baselines (Tab. 1, rows 3-8; Tab. 2), image-only RvTC achieves state-of-the-art performance on AVA with SRCC of 0.833 and PLCC of 0.831, surpassing the previous best method One-Align by 1.0 and 1.2 correlation points respectively. On IQA datasets, RvTC matches or exceeds existing state-of-the-art across all benchmarks: achieving equivalent performance to One-Align on KonIQ-10k and SPAQ, while substantially outperforming all methods on KADID-10k.
Table 3: The importance of task-level prompts and the model’s vocabulary. Results on AVA.
Method SRCC PLCC
Q-Align 0.822 0.817
[…]
Reversed Syllabus 0.8211 0.8180
Alternative Syllabus 11footnotemark: 1 0.8214 0.8193
Image-Only 0.8229 0.8197
[…]
RvTC - Image-Only (51 bins) 0.8329 0.8314
Ineffectiveness of current multimodal strategies. Tab. 3 reveals a key finding that challenges current assumptions about MLLM fine-tuning for regression. Neither human-based vocabulary constraints nor generic task-level prompts improve performance over image-only training. Fine-tuning mPLUG-Owl2 with its original vocabulary classification head and generic prompts yields nearly identical performance to our image-only RvTC model with 5 bins, the number of words used in the syllabus of Q-Align. This confirms that current multimodal approaches fail to leverage cross-modal understanding.
Impact of regression formulation. The progression from vocabulary-constrained classification to unconstrained bin-classification shows clear benefits. Moving from Q-Align’s 5-token vocabulary approach to RvTC with 5 bins yields similar performance (SRCC 0.823), but increasing to 51 bins provides a substantial boost to 0.833. This demonstrates that discretization granularity through increased bin count is more effective than attempting to align outputs with human vocabulary patterns, eliminating the need for manual vocabulary crafting while achieving superior performance.
Implications for multimodal approaches. These findings establish that current multimodal fine-tuning strategies provide no benefit over carefully designed image-only training. Fig. 3’s bin analysis shows that our framework offers a simple yet powerful alternative to complex distributional modeling approaches. However, this raises the question of whether multimodal capabilities provide additional value, a question we address in Sec. 4.3.
4.3 Impact of data-specific prompts
In this section, we investigate whether semantically meaningful prompts can unlock the cross-modal capabilities of MLLMs. We explore this by fine-tuning RvTC on the AVA dataset while incorporating challenge titles as data-specific prompts for each image.
Challenge titles as semantic descriptors. The AVA dataset contains rich semantic information in the form of challenge titles that characterize images with descriptive phrases. Examples include ”Rule of Thirds”, ”School Days Geometry”, ”Shoes” and ”Stationary”. These titles are concise yet semantically relevant, providing meaningful context that relates directly to the visual content and aesthetic properties being assessed.
Performance improvements. Tab. 1 demonstrates the substantial impact of incorporating challenge titles. In the linear probing setting (RvTC-LP+), adding challenge titles improves performance from 0.709 to 0.742 (average SRCC and PLCC), representing a significant gain of 3.3 correlation points without any backbone fine-tuning. This improvement demonstrates that language-based features enhance performance even when the multimodal backbone remains frozen.
The gains become even more pronounced with full fine-tuning. RvTC+ achieves a remarkable correlation of 0.90, improving from our already state-of-the-art image-only baseline of 0.83. This represents a jump of 7 correlation points and establishes a new state-of-the-art on the AVA dataset.
4.4 Ablation studies and analysis
This section provides ablation studies and analysis to understand the mechanisms behind RvTC’s performance improvements. We establish that a substantial portion of the gains achieved by incorporating challenge titles into RvTC fine-tuning on the AVA dataset stem from cross-modal understanding rather than statistical artifacts. Through controlled experiments, we decompose performance improvements into inter-challenge (driven by statistical biases) and intra-challenge (indicating semantic understanding) components. Additionally, we analyze the impact of bin count on performance across different training configurations, demonstrating that our framework achieves optimal performance through straightforward bin increase. We examine how data-specific prompts interact with our bin-based regression approach, showing that semantic prompts provide stabilizing effects that mitigate overfitting while maintaining performance gains.
Decomposing performance improvements. We distinguish between two types of performance improvements: Inter-challenge improvement refers to gains driven by differences in the statistical properties of challenge-specific data subsets. These improvements are reflected in per-challenge average predicted Mean Opinion Scores (MOS) and could potentially be achieved through statistical pattern recognition without semantic understanding. Intra-challenge improvement, conversely, is measured by correlation metrics of predictions within individual challenges. This type of improvement cannot be directly explained by per-challenge statistical biases and indicates the model’s ability to capture meaningful cross-modal relationships between textual and visual features.
Table 4: Ablation study analyzing whether improvements of RvTC on AVA using challenge titles stem from inter-challenge statistical biases by removing the semantic information from the challenge title during fine-tuning
Method SRCC PLCC
RvTC (image-only) 0.833 0.831
RvTC - Challenge ID 0.851 0.843
RvTC - Shuffled Titles 0.860 0.851
RvTC+ (with challenge titles) 0.899 0.901
Figure 2: Performance analysis of RvTC on AVA, comparing image-only RvTC (red) and incorporating image titles RvTC+ (blue) per-challenge average MOS predictions (top figure) and intra-challenge correlation of predictions (bottom figure) implying that the model is leveraging cross-modal features
Experimental design. To isolate these two sources of improvement, we conduct controlled ablation studies comparing image-only RvTC with RvTC+ across different prompt configurations (Tab. 4 and Fig. 2; see Tab. 5 for comparable analysis on AGIQA-3k).
In Tab. 4, ”RvTC - Challenge ID” we replace each challenge title in the textual prompt with a unique challenge identifier (positive integer string). This configuration allows the model to leverage inter-challenge statistical bias while eliminating semantic content.
In ”RvTC - Shuffled Titles” we randomly reassign challenge titles such that each challenge corresponds to a different title throughout training and evaluation. We ensure no challenge retains its original title. This disrupts semantic coherence while preserving the grouping effect that enables statistical bias exploitation.
In ”RvTC+ (with challenge titles)” we provide the full textual challenge title enabling exploitation of both statistical biases and rich semantic cues.
Key findings. Tab. 4 reveals that the overall improvement from incorporating challenge titles cannot be explained merely by inter-challenge statistical biases. The progression from image-only RvTC to full challenge titles demonstrates substantial gains beyond what statistical artifacts can account for.
Fig. 2 complements these findings. The analysis compares image-only RvTC (red) and RvTC+ (blue) performance within individual challenges containing at least 30 test images. The results demonstrate: (1) Enhanced per-challenge average MOS predictions (top): Improvements that can be partially attributed to inter-challenge statistical bias exploitation. (2) Improved intra-challenge correlation of predictions (bottom): Gains that indicate cross-modal feature utilization.
The substantial improvement in intra-challenge correlations provides compelling evidence that the model leverages semantic understanding rather than merely exploiting statistical patterns.
Implications. These findings establish that incorporating semantically meaningful challenge titles enables RvTC to access and utilize cross-modal understanding capabilities that are not utilized when training with generic prompts or image-only inputs. The decomposition analysis confirms that a significant portion of the observed performance gains stems from multimodal reasoning rather than statistical bias exploitation, validating the importance of semantic coherence in multimodal regression tasks.
Figure 3: Performance of image-only RvTC (top) and RvTC+ with challenge titles (bottom) on AVA using different number of bins and training lengths
Bin count analysis. Fig. 3 (top) demonstrates that performance scales monotonically with bin count across different training lengths, rendering the quantization scheme practically irrelevant once a sufficient number of bins is used. We systematically evaluated 5–101 bins, finding that performance improvement saturates at approximately 51 bins across all datasets. Notably, increasing the number of bins consistently improves correlation scores for all training durations, while longer training has a negative effect, especially for smaller bin counts, due to overconfidence in binning classification. This analysis reveals that optimizing the number of bins in RvTC is straightforward; unlike vocabulary-constrained approaches, our method achieves improved performance and training stability simply by increasing the randomly initialized regression head’s granularity.
Fig. 3 (bottom) reveals how data-specific prompts interact with our framework. Consistent with the image-only setting, increasing the number of bins consistently improves correlation scores across all training durations. However, challenge titles introduce a stabilizing effect: they completely eliminate the negative impact of prolonged training observed in the image-only case when the number of bins is sufficiently large. This suggests that incorporating challenge titles mitigates overfitting in scenarios with fewer bins and prolonged training.
4.5 Generalization to AI-generated images
To demonstrate the generalizability of our findings beyond aesthetic assessment, we evaluate RvTC on AI-generated image quality assessment using the AGIQA-3k dataset. This analysis serves two purposes: (1) validating that data-specific prompts improve performance across different domains, and (2) showing that semantic understanding, rather than statistical bias, drives the observed improvements.
Dataset and task formulation. AGIQA-3k contains image-prompt pairs with two evaluation tasks: alignment (how well the generated image matches the prompt instructions) and perceptual quality (visual quality of the generated image). Each image-prompt pair receives two mean opinion scores (MOS) corresponding to these tasks. The semantic content in prompts is directly relevant to the alignment task but less critical for perceptual quality assessment, providing an ideal testbed for examining when and how textual information contributes to regression performance.
Table 5: Performance of RvTC on AGIQA-3k’s Alignment and Perceptual tasks with different training and evaluation prompts. Bold indicates the best result on each task
Train Eval Alignment Task Perceptual Task
With prompt Image-Only Shuffled Prompt With prompt Image-Only Shuffled Prompt
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
With Prompt 0.810 0.889 0.687 0.826 0.634 0.702 0.872 0.916 0.868 0.901 0.872 0.911
Image-Only 0.715 0.817 0.692 0.826 0.679 0.756 0.869 0.905 0.878 0.918 0.865 0.884
Shuffled Prompt 0.672 0.828 0.678 0.827 0.676 0.828 0.872 0.914 0.862 0.902 0.872 0.915
Table 6: Prompt-Gated Regression on AGIQA-3k. Models fine-tuned on a single task (rows 1 and 2), compared to a model fine-tuned on both tasks (row 3). All models trained with textual prompts. Parenthesis shows difference from baseline.
Train Eval Alignment Task Perception Task
SRCC PLCC SRCC PLCC
Alignment 0.810 0.889 0.709 (-0.163) 0.793 (-0.123)
Perception 0.676 (-0.134) 0.781 (-0.108) 0.872 0.916
Alignment + Perception 0.804(-0.06) 0.885 (-0.04) 0.875 (+0.03) 0.913 (-0.03)
Experimental design for bias analysis. Tab. 5 systematically compares all training and evaluation combinations across both tasks to isolate semantic understanding from statistical artifacts. We evaluate three prompt configurations: (1) original prompts that maintain semantic coherence, (2) shuffled prompts where each image is randomly paired with a different prompt, disrupting semantic alignment while preserving statistical patterns, and (3) image-only that removes textual input entirely.
Evidence for semantic understanding. For the alignment task, training and evaluation with original prompts achieves optimal performance. Importantly, when a prompt-trained model is evaluated with shuffled prompts, performance degrades substantially below even the image-only baseline, demonstrating that the model has learned semantic associations rather than statistical patterns.
Task-specific prompt sensitivity. In contrast, the perceptual quality task shows minimal sensitivity to prompt variations. Performance remains largely stable across all prompt conditions (SRCC: 0.872$\pm$0.006), indicating that visual quality assessment relies primarily on image features rather than semantic context. This differential sensitivity validates that prompt utility depends on task semantics: prompts enhance performance when semantically relevant but provide little benefit for purely visual assessment tasks.
Note that when fine-tuning with shuffled prompts, both tasks produce an image-only model that largely ignores input prompts. This confirms that the model learns to disregard textual input when it lacks semantic coherence with the visual content.
Multi-task learning through prompt-gated regression. Tab. 6 demonstrates multi-task learning capabilities through what we term prompt-gated regression. By adding task identifiers (”Task: image alignment” or ”Task: image perceptual quality”) to prompts, a single model can simultaneously learn both regression tasks on the same data. The unified model achieves performance within 0.06 SRCC points of task-specific models, demonstrating that prompt formatting can effectively gate different regression objectives within a single framework. This extends previous results in [20] where it was shown that a single regression model can be trained on several datasets simultaneously.
Figure 4: Performance of RvTC fine-tuned on AGIQA-3k when evaluated with original prompt and with alternative prompt
Robustness to semantic paraphrasing. To further validate semantic understanding, we test whether the model can generalize to paraphrased prompts that preserve meaning while altering surface form. Using GPT-generated alternative prompts that rephrase original instructions with different structure and vocabulary while maintaining semantic content (Fig. 4), we evaluate whether performance depends on exact phrasing or underlying semantics.
Tab. 7 shows that performance remains stable when using semantically equivalent but syntactically different prompts. For both tasks, correlation scores show minimal variation (alignment: SRCC from 0.810 to 0.809; perceptual: SRCC from 0.872 to 0.874), confirming that the model captures semantic content rather than memorizing specific phrasings. This robustness to paraphrasing provides additional evidence that improvements stem from cross-modal understanding.
Implications for multimodal regression. These findings establish three key principles for effective multimodal regression: (1) prompt utility is task-dependent and tied to semantic relevance, (2) statistical bias cannot explain the observed improvements, as evidenced by performance degradation under semantic misalignment, and (3) models can achieve robust semantic understanding that generalizes across different linguistic expressions of the same concepts. Together, these results demonstrate that our approach successfully unlocks cross-modal capabilities in MLLMs for regression tasks when provided with semantically meaningful textual context.
Table 7: Performance of RvTC fine-tuned on AGIQA-3k with original prompts and evaluated with both original and alternative prompts
Prompt Task Alignment Task Perception Task
SRCC PLCC SRCC PLCC
Original Prompts
0.810 0.889 0.872 0.916
Alternative Prompts
0.809 0.889 0.874 0.917
5 Conclusions
This work challenges how multimodal large language models should be applied to image-based regression tasks. Our analysis reveals that effective multimodal regression requires moving beyond human-mimicking approaches. Three key insights emerge from our study: First, vocabulary constraints hinder rather than help regression performance; simple bin-based classification outperforms complex vocabulary-dependent methods. Second, semantic relevance in prompts is crucial for unlocking cross-modal capabilities, while generic task descriptions provide no meaningful benefit over image-only training. Third, architectural simplicity often trumps complexity; i.e. increasing classification bins vs. distributional modeling approaches. While our experiments focus on image assessment domains, the generalizability of these principles to other multimodal regression tasks remains an important area for future investigation. This work provides a foundation for developing more effective multimodal regression systems and underscores the critical role of semantic coherence in cross-modal understanding.
This work opens several research directions. First, investigating whether image-only fine-tuning protocols can preserve multimodal prompt-leveraging capabilities would enable more flexible and resource-efficient model development. Second, exploring multitask setups where regression and text generation tasks complement each other could lead to more capable unified models.